Thread View: sci.astro.hubble
1 messages
1 total messages
Started by hst-status
Mon, 28 Dec 1992 18:14
HST STATUS REPORT, 12/21 - 12/22
Author: hst-status
Date: Mon, 28 Dec 1992 18:14
Date: Mon, 28 Dec 1992 18:14
96 lines
4007 bytes
4007 bytes
HUBBLE SPACE TELESCOPE
DAILY REPORT #789
PERIOD COVERED: 0800 EST 12/21/92 - 0800 EST 12/22/92
Daily Status Report as of 357/1200Z.
1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED:
1.1 Completed FGS 4031 (A Search for Planetary Companions to Low Mass)
The Fine Guidance Sensors were used to measure the position of star
Proxima Centauri relative to background reference stars. The intent
is to determine the presence of a Jupiter Sized planet in orbit
around the target star. The observations were monitored by the OSS,
and no problems were reported.
1.2 Completed Three Sets of PC 4714 (Internal Delta Flats)
The Planetary Camera was used to take three different sets of
internal delta flat images. All of the observations were received
and displayed in the OSS. They all were reported to be okay.
1.3 Completed GHRS/2 3964 (Dynamics and Energy Balance in Stellar
Regions)
The High Resolution Spectrograph was used to observe the late type
star HD61421 which is known to have plasma heated to temperature
above 10,000 K above the photosphere. Four spectral observations
were received and displayed. A wave length calibration image was
also taken. A observations were reported to be okay.
1.4 Completed FOC/96 3428 (High Voltage Turn On)
A standard high voltage turn on test was performed using the Faint
Object Camera. Two internal flat, and one dark observations were
taken. This was strictly an internal test. All of the observations
were received and displayed. They all looked okay.
1.5 Initiated GHRS/2 2536 (Deuterium in the Local Interstellar Gas)
The High Resolution Spectrograph's Echelle spectrograph was used to
observe the interstellar absorption lines in the spectra of white
dwarf star G191-B2B. A total of five observations have been
completed , and a sixth observation has been initiated. The first
two observations included two deflection calibrations, and spiral
search, and peakup. the fifth observation was a calibration
spectrum. The sixth observation is the spectral observation of the
target star.
2.0 FLIGHT OPERATIONS SUMMARY:
2.1 Guide Star Acquisitions:
Scheduled Acquisitions : 3
Successful : 3
Scheduled Reacquisitions : 9
Successful : 9
Losses of Lock : 1
(LOL at 356/1312Z)
2.2 FHST Updates:
Scheduled : 5
Successful : 5
2.3 Significant HSTARs: None
2.4 Operations Note: None
3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD:
3.1 Complete GHRS/2 2536 (Deuterium in the Local Interstellar Gas)
3.2 Complete Two Sets of PC 4714 (Internal Delta Flats)
3.3 Complete Two Sets of S/C 4215 (FGS ITS Test)
3.4 Uplink Updated Ephemeris Tables, Per MOSES 2871
3.5 Complete FOS/RD 3816 (White Dwarf Stars)
3.6 Complete PC 4106 (HST Medium Deep Survey: Low Latitude)
3.7 Complete WFC-PC 4109 (Rapid Internal Monitor)
4.0 ANOMALY STATUS:
4.1 Significant New Anomalies: None
4.2 Significant Unresolved Anomalies: None
5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time
-END-
Thread Navigation
This is a paginated view of messages in the thread with full content displayed inline.
Messages are displayed in chronological order, with the original post highlighted in green.
Use pagination controls to navigate through all messages in large threads.
Back to All Threads